How many moles of silver atoms are in 1.8 * 10^20 atoms of silver?
1
Expert's answer
2012-12-11T08:38:41-0500
N - number of atoms
Na- Avogadro Constant
n-amount of element (compound) with some Molecular weight ( Mw )

n=N*Mw / Na.

Using this formula or trivial equation below
&
n 1.8x10^20
& Ag------->Ag
108 & 6.02x10^-23

n=108*1.8^20 / 6.02x10^23=0.0323 moles